# Context Matters: Pushing the Boundaries of Open-Ended Answer Generation with Graph-Structured Knowledge Context

### v2

@@ 5 Contextual Content Crux
7 Human Evaluation
In the results of the automatic evaluation, it is essential to remember that low values of BERTScore or FactSumm might also correspond to lower-quality ground truth answers posted by humans. Thus there are possibilities of the model generated answers to be superior in quality compared to the ground truth answers. Such points can be verified only by human judgement experiments which is what we conduct in this section.
8 Error Analysis
To better understand the quality of auto-generated answers, we examine randomly sampled answers and observe that they sometimes have errors originating from either the retrieval or generation phase, or both. This section presents a systematic error analysis highlighting the error types and corresponding examples (see Table 4).
9 Conclusion
Driven by these encouraging results, we aim to develop GraphContextGen into an accessible tool. We envision a straightforward task, perfect for novice developers, users, and researchers, where they can obtain the optimal answers from GraphContextGen. This initiative highlights the significant potential of state-of-the-art generative language models in bridging knowledge grounding gaps and enhancing factual coherence, especially given hardware constraints. We look forward to more advancements in this area.
Appendix A Appendix
A.1 Instruction tune hyperparameters
Hyperparameter Value
Learning Rate
2e-4
Batch Size
4
Gradient Accumulation Step
1
Number of Epochs
10
Weight Decay
0.001
Optimizer
paged adamw 32bits
LR scheduler
cosine
Warmup ratio
0.03
Max grad norm
0.3
bf16
True
LoRA r, alpha, dropout
64, 16, 0.1
bf16
True
Quantization
4bit
PEFT Techniques
LoRA
Trainer
SFTT
A.2 Pre-LLM era baselines
A.2.1 AnswerBot Xu et al. [2017], Cai et al. [2019]:
A.2.2 GenQA Hsu et al. [2021]:
A.2.3  TechSumBot Yang et al. [2023a]:
A.3 Ablation study
  1. (i)
    Based on embedding algorithms: The performance differences between sentence BERT Reimers and Gurevych [2019] and BGE embeddings Xiao et al. [2023] within the TextGen model are evident. The BGE variant demonstrates a slightly improved performance over sentence BERT.
  2. (ii)
    Based on one-shot: When focusing on one-shot methodologies, the GraphGen model achieves marginally better scores compared to the TextGen model. This suggests that the incorporation of graph structures potentially aids in better knowledge retrieval.
  3. (iii)
Method AskUbuntu Unix ServerFault
Based on embedding algorithms
[w/o INST] TextGen (Sentence BERT)
0.808 0.783 0.803
[w/o INST] TextGen (BGE)
0.812 0.809 0.810
Based on one-shot
[w/o INST] TextGen
0.809 0.807 0.807
[w/o INST] GraphGen
0.818 0.809 0.812
Based on only knowledge graph context
[w/o INST] OnlyContextGen[FT]
0.819 0.807 0.819
[w/o INST] OnlyContextGen[FG]
0.827 0.820 0.818
A.4 Sample prompt and generated answer

### v3

@@ 5 Contextual Content Crux
A.2 Pre-LLM era baselines
A.2.1 AnswerBot Xu et al. (2017); Cai et al. (2019):
A.2.2 GenQA Hsu et al. (2021):
A.2.3  TechSumBot Yang et al. (2023a):
A.3 Sample prompt and generated answer
A.4 Sample questions retrieved from our method
We include a few examples in Table 6 that take into account both PPR (graph structure-based) and simple similarity for certain questions. The questions retrieved by the PPR method are very specific, to-the-point, and strongly related to the actual question. The simple similarity-based questions, on the other hand, are very generic (e.g., What are some common network troubleshooting tools…, What are the best practices for managing log files…).
A.5 Enhanced context formulation
Figure 4: Sample context preparation.
A.6 Human annotation